Two tickets currently open:
683120, backoff limit failures: we have converged on a fix, which requires that the pod storage limit be exposed as an environmental variable so that the pilot can use that to determine the available space. Fernando has added the necessary environment variables, now we are waiting on Paul.
1000970, tape transfers failing: we will put the tape into a writing downtime to allow enough space to clear up for the transfers to resume.
Some job failures due to ongoing dcache/network upgrades and SENSE tests. Over the past week these have been eliminated thanks to the (temporary) addition of two powerful servers borrowed from NESE, which have made rebalancing and draining pools for upgrades go much more smoothly.
800 Gbps connection to NESE now available, next step is to get 800 Gbps to ESNET.